OpenStudy (anonymous):

how do you graph this linear equation y=3x+4

OpenStudy (anonymous):

if y = 0: 0=3x+4 x=-4/3 when x=0, y=4 in ur x-axis, find the point where x = -4/3 in ur y axis find the point where y= 4 connect both of the line , that is the graph, u can extend it to wherever u want

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Get some points by trying out values for x, and seeing what value of y they have: x 0 1 2 3 3x+4 3(0)+4 3(1)+4 3(2)+4 3(3)+4 y 4 7 10 13 Draw them!